OpenStudy (brainzonly):

As they say, beggars can't be choosers, in fact begger take what they can get. A begger on the street can make one cigarette out of every 6 cigarette butts he finds. After one whole day of searching and checking public ashtrays the begger finds a total of 72 cigarette butts. How many cigarettes can he make and smoke from the butts he found?

OpenStudy (mathmale):

One way of solving this problem would be to use ratios. 6 butts 72 butts ----------- = -------------- 1 cigarette ? cigarettes

OpenStudy (brainzonly):

Ah huh. My internet is sooooooooooooo slow I might not respond for a while. (I should name my internet connection SnailNet XD)

OpenStudy (brainzonly):

I think it's \[6=\frac{ 72 }{ x }\]

OpenStudy (mathmale):

Explain how you could solve for x and then check that result.

OpenStudy (irishboy123):

that's not quite right IMHO he make make some more smokes from the x smokes he makes from the orig 72 butts

OpenStudy (brainzonly):

yay, snailnet finally got me here!!! anyway... multiply 72 to both sides.. 400 something O.O No way.

OpenStudy (kevin):

So it will become \[\frac{ 6 }{ 1 } = \frac{ 72 }{ x }\] then \[x = \frac{ 72 \times 1 }{ 6 }\]

OpenStudy (mathmale):

Can't \[x = \frac{ 72 \times 1 }{ 6 }\]be simplified?

OpenStudy (brainzonly):

O MY GOD!!! I'M SO STUPID!!!!!!!!!!! Look at it again. THIS PROBLEM IS TRICKING YOUR BRAIN INTO THINKING THAT THE PROCESS IS MORE COMPLICATED THAN IT IS!!!!!! If the begger can make a whole cigarette from 6 butts then he can make 12 cigarettes from the 72 he found. Once he smokes those, he then will have another 12 butts, which gives him enough to make another 2 cigarettes. A total of 14. LOL!!!!

OpenStudy (brainzonly):

Party over XD I love brain teasers!!! :DDDD

OpenStudy (kevin):

you can divide the numerator with the denominator :D

OpenStudy (brainzonly):

72/6 equals 12...

OpenStudy (kevin):

Party over..... :o Oh my god :(

OpenStudy (kevin):

That's it :D 12

OpenStudy (brainzonly):

and then he makes 2 more smokes out of that 12/6=2 so 12+2 equals 14 :DDDDDD

OpenStudy (kevin):

Then he will get cancer :D

OpenStudy (brainzonly):

@Kevin WRONG!!!! This entire problem is gauged to trick you XD!!!!!!!!!!!!!!!!! cancer LOL

OpenStudy (brainzonly):

ready for the next 1? this one's even better :DD

OpenStudy (brainzonly):

?

OpenStudy (kevin):

Alright, that will be the last one because I'm going to sleep after this :D

OpenStudy (kevin):

Never mind my word before :p